40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Sep 26 19:48:09 2016, Lydia, Update, SUS, ITMX stuck again, ITMY whitening issue 44.png26.png
    Reply  Tue Sep 27 08:49:47 2016, Steve, Update, SUS, seismic activity is up 4.3mSaltonSee.pngITMXstuck.png
       Reply  Mon Oct 10 09:31:25 2016, Steve, Update, SUS, PRM damping restored 3.7mLomaLinda.pngETMY_UL_glitch.pngMC_glitcing?.png
       Reply  Tue Nov 1 09:03:08 2016, Steve, Update, SUS, seismic activity is up seismicActivity.png
          Reply  Thu Dec 8 07:56:05 2016, Steve, Update, PEM, EQ6.5m Northen CA eq6.5FerndaleCA.png16d_glitching-_trend.pngEQ6.5_&4.7mFerndaleCa.png
             Reply  Thu Dec 29 10:24:56 2016, Steve, Update, PEM, EQ5.7mHawthorn NV eq5.7HawthorneNV.png
    Reply  Tue Sep 27 18:04:50 2016, Lydia, Update, SUS, ITMX slow channels down, ITMY diagonalization update ITMY_diagsuccess.pdf
Message ID: 12520     Entry time: Tue Sep 27 18:04:50 2016     In reply to: 12518
Author: Lydia 
Type: Update 
Category: SUS 
Subject: ITMX slow channels down, ITMY diagonalization update 

[Teng, Lydia]

When we plugged in the back cables yesterday on the whitening boxes after switching them, two of the ITMX PDMon channels (UR and LR) got stuck at 0. This caused me to believe ITMX was still stuck even when it was freed. However, it was left in a stuck state overnight and freed again today after this issue was discovered. The alignment sliders have been set to 0 as a safety net to keep ITMX from getting stuck again if c1susaux is restarted again. We switched the cables back and the problem was still there.

The ITMY UL whitening filter problem, which the cables were originally switched to diagnose, was also still there. Ericq suggested we turn off all the whitening filters in order to get diagonalization data that would not show a phase difference between coils. We ran the diagonalization again with all the dewhitening filters off and got much cleaner results, with no visible cross-coupling peaks remaining between the degrees of freedom (see attachemnt 1). We did not apply this matrix to the damping, however, because there are elements which have the wrong sign compared to the ideal matrix. Significant adjustments to the output matrix will probably need to be made if this result is to be used. We also verified that the phase problem had been solved in DTT, where we saw the same sign discrepancies as in the matrix below. 

Damping can be turned back on, using the old, non-diagonalized matrix currently in effect. There is enough free swing data to diagonalize ITMY now, so feel free to mess with it. 

Matrix (wrong signs red, suspiciously small elements orange):

           pit     yaw     pos         side    butt
UL    1.633   0.138   1.224   0.136   0.984  
UR   -0.202  -1.768   1.179   0.132  -1.028  
LR   -2.000   0.094   0.776   0.107   1.001  
LL   -0.165   2.000   0.821   0.111  -0.987  
SD    0.900   1.131  -1.708   1.000  -0.107  

 

Attachment 1: ITMY_diagsuccess.pdf  32 kB  | Hide | Hide all
ITMY_diagsuccess.pdf
ELOG V3.1.3-